a person can see the top of a building at an angle of 65 degrees the person is standing 50 ft away from the building and has an eye level of 5 ft how tall is the building to the nearest tenth of a foot
Answer by lwsshak3(11628)

a person can see the top of a building at an angle of 65 degrees the person is standing 50 ft away from the building and has an eye level of 5 ft how tall is the building to the nearest tenth of a foot.
***
let x= distance from eye level to top of bldg.
x/50=tan 65˚
x=tan 65˚*50
Height of building=x+5=tan 65˚*50+5=112.2 ft